How do I find the equation of the line perpendicular to 5y=x+3 that passes through the point (8,6)
And how do I write it in slope intercept form?

Answer by Alan3354(69443)

How do I find the equation of the line perpendicular to 5y=x+3 that passes through the point (8,6)
And how do I write it in slope intercept form?
-----------------
Find the slope of the given line, m1.
Then, the slope of lines perpendicular is the negative reciprocal, m2 = -1/m1.
------
Thru the point with slope m ---> y-y1 = m*(x-x1) where (x1,y1) is the point.